How to change in total surplus. consumer and producer surplus?
Old equilibrium point is Q=5064 P=0.22
the new equilibrium point is Q=3984 P=0.27.

Many thanks!


supply 1.60 New supply (P) 1.20 1.19 1.08 1.00 0.98 0.89 0.80 ◇ 0.71 8 0.40 .47 0.20 0.00 Quantity, Q (thousands) prices and quantities 在这里输入你要搜索的内安
n Suppl tu (c) From your chart, what can you say about the change in total surplus, consumer surplus, and producer surplus as a result of the supply shock? (Hint: You may find the following information useful: the old equilibrium point is Q 5,064, P 0.22; the new equilibrium point is Q- 3,984, P 0.27 Part 7.2 Interpreting supply and demand curves
Answer #1

Consumer surplus is the area under demand curve that represents the difference between the consumer is willing to pay and how much he actual pays. Producer surplus is the area under supply curve that represents the difference between what a producer is willing able to accept for selling a product and what the producer sells it for. The area of these can be calculated as 1/2*BASE*HEIGHT (area of triangle). Refer to this diagram.

10 Thursday |ONum wws odRefer to the situation given in the question, Friday 2t 12 Saturday

Total surplus(TS)=Consumer Surplus(CS)+Producer Surplus(PS)

Initial Consumer surplus=1/2*5064*(1.6-.22)=3494.16

Initial producer surplus=1/2*5064*(.22)=557.04

Initial total surplus=4057.2

New consumer surplus=1/2*3984*(1.6-.27)=2649.36

new producer surplus=1/2*3984*.27=537.84

Total new surplus=3187.2

CS have decreased, PS have decreased so TS have also decreased